OpenStudy (anonymous):

A disk rotates with constant angular acceleration. The initial angular speed of the disk is 2π rad/s. After the disk rotates through 15π radians, the angular speed is 13π rad/s. a) What is the magnitude of the angular acceleration? (b) How much time did it take for the disk to rotate through 15π radians? (c) What is the tangential acceleration of a point located at a distance of 8 cm from the center of the disk?

OpenStudy (roadjester):

\[\large \omega_i = 2\pi {\frac {rad} s} \] \[\large \omega_f = 13\pi {\frac {rad} s} \] \[\large \theta_f = 15\pi { {rad} } \] \[\large \omega_f=\omega_i+\alpha t\]

OpenStudy (roadjester):

@vegeto no time was given?

OpenStudy (roadjester):

Ok, it looks like the first thing you'll need to do is solve for the angular acceleration. \[\huge \omega_f^2=\omega_i^2+2\alpha(\theta_f-\theta_i)\] Once you have the angular acceleration, you can find the time. Tangential acceleration is simply \[\huge a=r\alpha\] Just make sure your units make sense.